Your teacher is giving you a test worth 100 points containing 40 questions. There are two-point and four-point questions on the test. How mamy of each type of question are on the test? please help me to sovle it. thank you.

Answer by checkley71(8403)

2X+4(40-X)=100 OR 2X+160-4X=100 OR -2X=-60 OR X=30 2 POINT QUSTIONS & 40-30=10
4 POINT QUESTIONS
PROOF 30*2+10*4=60+40=100 OR 100=100